I have 7 bags of marbles. There are 200 marbles in total. Using the clues below, work out how many marbles are in each bag.
Bag 1 + Bag 2 = 57 marbles
Bag 2 + Bag 3 = 83 marbles
Bag 3 + Bag 4 = 71 marbles
Bag 4 + Bag 5 = 43 marbles
Bag 5 + Bag 6 = 66 marbles
Bag 6 + Bag 7 = 43 marbles
Thank you so much for your help. I tried to figure this out but came to no avail.

Found 2 solutions by jim_thompson5910, rapaljer:
Answer by jim_thompson5910(28595)

Lets use these variables for each bag:
Bag 1 = a
Bag 2 = b
Bag 3 = c
Bag 4 = d
Bag 5 = e
Bag 6 = f
Bag 7 = g

Now set up the following system of equations

a + b = 57 
    b + c = 83 
        c + d = 71 
            d + e = 43
                e + f = 66
                    f + g = 43 



Also since we know how many marbles there are total, we can use this equation:

a+b+c+d+e+f+g=200



Now subtract (a+b=57), (c+d=71), and (f+g=43) from a+b+c+d+e+f+g=200 to eliminate everything but one bag (in this case bag 5 which is denoted "e")

  a+b+c+d+e+f+g=200
-(a+b          =57)
-(    c+d      =71)
-(          f+g=43)
-------------------
          e    =29


So we know that bag 5 has 29 marbles

Now plug in e=29 to find f

Bag 6:

29 + f = 66

f = 37


So we know that bag 6 has 37 marbles
Now plug in f=37 to find g

Bag 7:

37 + g = 43

g = 6



So we know that bag 7 has 6 marbles
Now plug in e=29 to find d

Bag 4:

d + 29 = 43

d = 14



So we know that bag 4 has 14 marbles
Now plug in d=14 to find c

Bag 3:

c + 14 = 71

c = 57


So we know that bag 3 has 57 marbles
Now plug in c=57 to find b

Bag 2:

b + 57 = 83

b = 26



So we know that bag 2 has 26 marbles
Now plug in b=26 to find a

Bag 1:

a + 26 = 57

a = 31


So we know that bag 1 has 26 marbles

So here's a summary of all of the bags:
Bag 1 = 31 marbles
Bag 2 = 26 marbles
Bag 3 = 57 marbles
Bag 4 = 14 marbles
Bag 5 = 29 marbles
Bag 6 = 37 marbles
Bag 7 = 6 marbles

Check:
31+26+57+14+29+37+6=200
200=200

Answer by rapaljer(4667)

What you have here is 7 equations and 7 unknowns, unless someone sees a faster way to do this!!

Let the 7 variables be x1, x2, x3, x4, x5, x6, and x7 respectively.

Now, here are the equations:
x1+ x2+ x3+ x4+ x5+ x6+x7=200
x1+x2=57
x2+x3=83
x3+x4=71
x4+x5=43
x5+x6=66
x6+x7=43

If you have a graphing calculator, like a TI85, 86, 83+, or 84, you may have a program that will solve this system called [SIMLT] or perhaps [POLYSMLT].

If you don't have a calculator, then try getting everything in terms of x1. Do this by starting with
x2 =57-x1

Then
x3 = 83-x2
x3=83-(57-x1)
x3=26+x1

Next,
x4= 71-x3
x4=71-(26+x1)
x4=45-x1

x5=43-x4
x5=43-(45-x1)
x5=-2+x1

x6=66-x5
x6=66-(-2+x1)
x6=68-x1

x7=43-x6
x7=43-(68-x1)
x7=-25+x1


Now,
x1+ x2+ x3+ x4+ x5+ x6+x7=200
x1+(57-x1)+(26+x1)+(45-x1)+(-2+x1)+(68-x1)+(-25+x1)=200
x1+169=200
x1=31
x2=26
x3=57
x4=14
x5=29
x6=37
x7=6

That should do it!!!

R^2 at SCC